a element has density of 20 g/cm^3. what is the radius of a sphere that has a mass of 70 kg of this material?
Answer #1

ρ=20g/cm3, m=70kg = 70000g

Use equation,

Density = mass/volume

ρ=m/V

V= m/ρ

(4/3*πr3) = m/ρ

r3= 3πm/4ρ

r= (3πm/4ρ)1/3

Plug values,

r = [(3*3.14*70000)/(4*20)]^1/3

r = 20.20 cm